OpenStudy (jenniferjuice):

The number of hours (H) that a candle will burn increases when the length of the candle (L) increases. Write the correct equation for this scenario, and solve for the number of hours when the length is 2. length hours 15 3 20 4 A. H = .2L; H = .4 B. H = 1L; H = 2 C. H = 45 over L; H = 22.5 D. H = 45 over L; H= 3

OpenStudy (anonymous):

You see, 15/20 = 3/4. So the relationship is linear. Let's take it as L = k*H, k is constant to be determined. Put L= 15 and H = 3. Find k. Now use this relation to find H for L = 2